OpenStudy (anonymous):

determine the coefficient of x^5 in the expression (2x-3)^8 help pls!!

OpenStudy (anonymous):

going with binomial theorem\[(a+b)^n=\sum_{k=0}^{n} \binom{n}{k} a^k b^{n-k}\]and setting a=2x ,b=-3 and n=8 ... will be your start point

OpenStudy (anonymous):

\[\sum_{k=0}^{8} \left(\begin{matrix}8 \\ 5\end{matrix}\right) 2^5 (-3) ^3 x^5\] Is this is?

OpenStudy (anonymous):

u just need to evaluate one term of the sum and its where k equals to 5 so just \[ \left(\begin{matrix}8 \\ 5\end{matrix}\right) 2^5 (-3) ^3 x^5\]